An electrical firm manufactures light bulbs that have a length of life that is normally distributed with mean equal to 778 hours and a standard deviation of 43 hours. Find the probability that a bulb burns between 755 and 892 hours.


1
Expert's answer
2021-03-02T05:11:09-0500

Let "X=" a length of life of the light bulb: "X\sim N(\mu, \sigma^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=778\ hr, \sigma=43\ hr"


"P(755<X<892)=P(X<892)-P(X\leq755)"

"=P(Z<\dfrac{892-778}{43})-P(Z\leq\dfrac{755-778}{43})"

"\approx P(Z<2.651163)-P(Z\leq-0.534884)"

"\approx0.9959892-0.2963651\approx0.699624"

The probability that a bulb burns between 755 and 892 hours is 0.699624.